what kind details
3. Recommend in years) for followi amount. a type of life insurance (including the dollar amount and length of the policy r the following people. Be sure to EXPLAIN how you came up with the SSS 7. She A) Sue is 26 years old, a single mom with two children, Jake age 5 and Grace- age currently makes $45,000 a year as a registered nurse. She has $9,000 in an emergency owns her 1994 Nissan Altima, has $28,000 in her companys 401K, and owns her house (market value of $89,000 with a mortgage of $54,000). She has no other debt. B) Joe is 26 years old and single. He has no dependents and makes $65,000 a year as a commodities trader. He rents a studio apartment for $1500 a month that is nicely furnished. He has no emergeney fund and no retirement or savings. He owns a really cool Porsche Boxter... (well not really owns it since the loan balance is $52,000). He has no other debt. C) Monica and Manuel Juarez are a dual-career couple who just had their first child. Manuel, age 29, already has a group life insurance policy for $200,000, but Monicas employer does not offer life insurance. An insurance agent is recommending that the 25-year-old Monica buy a S150,000 whole life policy that carries an annual premium of $1,370. They are currently renters, have $4,000 in an emergency fund, and own their two cars: a 2004 Nissan Altima and a 1995 Toyota T-100. They have no other debt and Manuel makes $45,000 a year and Monica makes $55,000. What do you think? EXPLAIN and make recommendations for life insurance for both of them.

media%2F3fd%2F3fd497a5-0ae6-44c0-bf0f-63
Answer #1

a) Sue age 26y, average retirement age is 60y, and has 2 children, average inflation rate is 2.3%, average monthly House hold expences is $1483.

by taking monthly taking as base

her monthly expenses = $1483*12 = $17796

and by consdering inflation the expenses can increase upto her retirement = FV(0.023,34,-17796,0,0)

= $902616  

and by adding her mortgage amount to the above amount =$902616+$54000 =$956616

so i suggested her to take $ 1000000 Term policy up to the age of 60years

b) by considering above procedure

i am suggeste him to take $ 400000 term policy upto age he get dependents because he don't have any dependents

c)i recommended both of them to take Term policies becaues the premium is very low as compared to treditional policies by using above proceses as by taking there annual income and inflation.

Note: the above answers are to the best my knowledge and these are different from there risk and life style and living standards and the inflation of the country.

2) a) Age of the property = 20y

age of the property at the time of destroyed = 15y

property has 15/20= 75% is 25% remaining life time

replacement value = $60000

ACV = $60000*25%= $15000-$1000

=$14000

b) = $60000-$1000 = $59000 (replacement cost)

Replacement cost means the claim is settled at the price that takes to reeplace that building so, here the replacement cost is $ 60000 minus deductable $1000 = $59000.